OpenStudy (carolina4567):

Need Help Asap Have A Few Other Questions Also ! Please Help Solve: -5x = 30 6 -6 25 35

OpenStudy (carolina4567):

I would say 6

OpenStudy (carolina4567):

Or -6 not sure

OpenStudy (eric_d):

It's -6

OpenStudy (carolina4567):

Thanks @eric_d

OpenStudy (carolina4567):

I got a few more like this one

OpenStudy (carolina4567):

Solve: -2x - 15 = 6x + 9 3 -3 -1 6

OpenStudy (anonymous):

-3

OpenStudy (eric_d):

Let me show you hw to get -6 @carolina4567

OpenStudy (carolina4567):

Thank u so much both And ok

OpenStudy (carolina4567):

-5x = 30 -5 / 5 = -6x

OpenStudy (carolina4567):

am i rght

OpenStudy (eric_d):

 @carolina4567